In a dielectric material having ε = 12 the displacement current is 25 times greater than
the conduction current at 100 MHz . The conductivity of the dielectric is
1
Expert's answer
2020-06-22T11:20:14-0400

Assume that we have a conduction current "J_c", displacement current "J_d".

The conductivity will be


"\sigma=\omega\epsilon\epsilon_0\frac{J_c}{J_d},\\\sigma=2\pi\cdot10^8\cdot12\cdot8.85\cdot10^{-12}\cdot\frac{1}{25}=0.00267\text{ S/m}."